OpenStudy (anonymous):

: Linear Algebra: Let B = (S^-1)AS, and x be an eigenvector of B belonging to an eigenvalue L. Show that Sx is an eigenvector of A belonging to L.

OpenStudy (anonymous):

The setup to this question indicates that \[Bx = Lx\] where L is a constant and B is the matrix \[B = (S^{-1}AS)\] Which means that \[A = SBS^{-1}\] It follows that \[A(Sx) = SBS^{-1}(Sx) = SB(x) = S(Lx) = L(Sx) \] So Sx is indeed an eigenvector of A with eigenvalue L.

OpenStudy (anonymous):

Thank you, that reminded me to use the fact that Bx = Lx, which I forget so easily in the middle of all these hw..

OpenStudy (anonymous):

No problem :)